Two cars of unequal masses use similar tyres.If they are moving with at the same initial speed,the minimum stopping distance(a)is smaller for the heavier car(b)is smaller for the lighter car(c)is same for both cars(d)depends on the volume of the car

Grade:Upto college level

1 Answers

pratham ashish
14 years ago
Hi, Ans to ur Qn is(c)the distance is same for both cars See how : the friction force is mue*m1*g (mue = friction coeff.) for other the friction force is mue*m2*g (mue is same for both as both have same type of tyres) but retardetion due to friction for I car = mue*m1*g/m1 = mue*g for II car = mue*m2*g/m2 = mue*g retardation is same for both , so time taken in stopping will also be same As s = ut+ (at^2)/2 HAVE NICE LUCK